OpenStudy (anonymous):

Johnny has been planting blueberries in a garden in his backyard. Johnny started with blueberries in one square foot of garden modeled by the function s(x) = 10. He is allowed to convert more of the garden, and each additional square foot he maintains allows his blueberry plants to produce more by a rate of a(x) = 1.4^(x-2). Explain to Johnny how to create an equation to predict the number of blueberries he can expect based on the number of square feet he maintains. Describe how to determine the number of blueberries he will grow with 5 square feet HELP!!!!

OpenStudy (anonymous):

@ganeshie8 @mathmale @misty1212 @thomas5267

OpenStudy (aaronandyson):

Row 1 = 15 Row 2 = 15 + 4 = 19 Row 3 = 15 + 4 + 4 = 23 Row 4 = 15 + 4 + 4 + 4 = 27 Row 5 = 15 + 4 + 4 + 4 + 4 = 31 Row n = 15 + 4(n - 1)